# Slender Trees and the Approximation Property

### v1

@@ 7. Making $\operatorname{ISP}$ indestructible under directed closed forcing
We can view $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)*\dot{\mathbb{L}}$ as an ordering on a product $\mathbb{P}\times\mathbb{Q}$ by letting $\mathbb{P}:=\operatorname{Add}(\tau,\kappa\smallsetminus\xi)$ and $\mathbb{Q}$ consist of tripels $(q,r,\sigma)$ with $q,r$ as in $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)$ and $\sigma$ such that $\Vdash\sigma\in\dot{\mathbb{L}}$. This ordering is easily seen to be iteration-like. Moreover, the base ordering is $\tau^{+}$-Knaster (this is clear) and the term ordering is ${<}\,\mu$-closed: Given a descending sequence $(p,(q_{\alpha},r_{\alpha},\sigma_{\alpha}))_{\alpha<\delta}$ for $\delta<\mu$, first find a lower bound $(p,(q,r))$ in $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)$. Then $(p,(q,r))$, being a lower bound, forces the sequence $(\sigma_{\alpha})_{\alpha<\delta}$ to be descending in the ${<}\,\kappa$-directed closed forcing $\dot{\mathbb{L}}$. Hence there exists a lower bound $\sigma$. In summary, $(p,(q,r,\sigma))$ is a lower bound of $(p,(q_{\alpha},r_{\alpha},\sigma_{\alpha}))_{\alpha<\delta}$. Now apply Theorem 3.7. ∎
Now we can show that $\mathbb{M}_{3}^{l}(\tau,\mu,\kappa)$ forces $\operatorname{ISP}$ in such a way that it is indestructible under ${<}\,\kappa$-directed closed forcing.
@@
Let $\tau<\mu<\kappa$ be regular cardinals such that $\tau^{<\tau}=\tau$ and $\kappa$ is supercompact. Let $l$ be a Laver function. Then $\mathbb{M}_{3}^{l}(\tau,\mu,\kappa)$ forces $\operatorname{ISP}(\tau^{+},\kappa,\geq\kappa)$ and it is indestructible under ${<}\,\kappa$-directed closed forcing.
@@
Let $\mathbb{M}(\nu):=\mathbb{M}_{3}^{l}(\tau,\mu,\nu)$. It suffices to show that for any $\mathbb{M}(\kappa)$-name $\dot{\mathbb{L}}$ for a ${<}\,\kappa$-directed closed forcing, $\mathbb{M}(\kappa)*\dot{\mathbb{L}}$ forces $\operatorname{ISP}(\tau^{+},\kappa,\geq\kappa)$. Additionally, it suffices to show that $\operatorname{ISP}(\tau^{+},\kappa,\lambda)$ is forced for arbitrarily large $\lambda$. Thus let $\dot{\mathbb{L}}$ be an $\mathbb{M}(\kappa)$-name for a ${<}\,\kappa$-directed closed forcing, let $\lambda\geq|\dot{\mathbb{L}}|$ and $\dot{f}$ an $\mathbb{M}(\kappa)*\dot{\mathbb{L}}$-name for a ${<}\,\tau^{+}$-slender $(\kappa,\lambda)$-list, forced by some $(p,\sigma)\in\mathbb{M}(\kappa)*\dot{\mathbb{L}}$. Let $\dot{F}$ be an $\mathbb{M}(\kappa)*\dot{\mathbb{L}}$-name for a function such that any $M\prec H(\Theta^{\prime})$ closed under $\dot{F}$ witnesses the slenderness. We will view $\mathbb{M}(\kappa)*\dot{\mathbb{L}}$ as a partial order on $\lambda$. Let $\Theta$ be large and $M\prec H(\Theta)$ with the following:
@@
Then $p$ is actually a condition in $\mathbb{M}(\nu)$ and $\pi(\sigma)$ is an $\mathbb{M}(\nu)$-name for a condition in $\pi(\dot{\mathbb{L}})$. Thus we can view $(p,\pi(\sigma))$ as a condition $q$ in $\mathbb{M}(\nu+1)$. Let $G_{0}$ be an $\mathbb{M}(\kappa)$-generic filter containing $q$ and work in $V[G_{0}]$. The collapse $\pi\colon M\to N$ extends to $M[G_{0}]\to N[\pi[G_{0}\cap M]]$ and, by the $\nu$-cc. of $\mathbb{M}(\nu)$, $G_{0}^{\prime}:=\pi[G_{0}\cap M]$ is $\mathbb{M}(\nu)$-generic over $V$. Let $H_{0}^{\prime}$ be the $\pi(\dot{\mathbb{L}})$-generic filter induced by $G$. Then $\pi^{-1}[H_{0}^{\prime}]\cup\{\sigma^{G}\}$ is (by elementarity) a ${<}\,\kappa$-sized collection of pairwise compatible conditions in $\dot{\mathbb{L}}^{G}$. Ergo there exists a lower bound $r$. If now $H_{0}$ is $\dot{\mathbb{L}}^{G}$-generic containing $r$, $\pi[H_{0}\cap M]=H_{0}^{\prime}$ is $\pi(\dot{\mathbb{L}})$-generic over $V[G_{0}^{\prime}]$. This shows that for any model $M$ as above, there exists a condition $q_{M}\leq(p,\sigma)$ forcing that for any $\mathbb{M}(\kappa)*\dot{\mathbb{L}}$-generic filter $G*H$, $\pi[G*H\cap M]$ is equal to the $\pi(\mathbb{M}(\kappa)*\dot{\mathbb{L}})$-generic filter induced by $G$. Now we transform, as before, $\dot{f}$ into a ground-model $(\kappa,\lambda)$-list. Let $a\in[\lambda]^{<\kappa}$.
@@
For simplicity we write $\pi_{M[G_{1}*H_{1}]}=\pi$. Toward a contradiction, assume that $\pi[\dot{f}(\check{a})]\notin V[G_{1}^{\prime}*H_{1}^{\prime}]$. As $V[G_{1}*H_{1}]$ is an extension of $V[G_{1}^{\prime}*H_{1}^{\prime}]$ by a forcing with the ${<}\,\tau^{+}$-approximation property, there exists $z\in[\pi(\lambda)]^{<\tau^{+}}\cap V[G_{1}^{\prime}*H_{1}^{\prime}]$ such that $z\cap\pi[\dot{f}^{G_{1}*H_{1}}(a)]\notin V[G_{1}^{\prime}*H_{1}^{\prime}]$. Because $H_{1}^{\prime}$ is generic for a ${<}\,\nu$-directed closed forcing, $z\in V[G_{1}^{\prime}]$ and by Lemma 4.6, $z\in N[G_{1}^{\prime}]$, so $z=\pi(y)$ for $y\in M[G_{1}]\subseteq M[G_{1}*H_{1}]$. Furthermore, $M[G_{1}*H_{1}]\cap H(\Theta^{\prime})$ is closed under $\dot{F}^{G_{1}*H_{1}}$, so it witnesses the ${<}\,\tau^{+}$-slenderness of $\dot{f}^{G_{1}*H_{1}}$. But then
@@
a contradiction. So $\pi[\dot{f}^{G_{1}*H_{1}}(a)]\in V[G_{1}^{\prime}*H_{1}^{\prime}]$ and there exists a condition $q_{a}$ as required. Lastly, we want to show that $q_{a}$ forces that $M[\Gamma]$ is a $\lambda$-ineffability witness for $\kappa$ with respect to $\dot{f}$. To this end, let $G_{2}*H_{2}$ be $\mathbb{M}(\kappa)*\dot{\mathbb{L}}$-generic containing $q_{a}$. By assumption we are in case (1) and there is $b_{e}\in M$ such that $b_{e}\cap M=e(M\cap\lambda)$. Define
@@
where the last equality holds because $q_{a}\leq q_{M}$, which implies $M[G_{2}*H_{2}]\cap V=M$. Let $\beta\in b_{f}\cap M$. By elementarity there is $\alpha\in G_{2}*H_{2}\cap M$ such that $\langle\alpha,\beta\rangle\in b_{e}$ and we have $\langle\alpha,\beta\rangle\in b_{e}\cap M=e(a)=\pi^{-1}(g(a))$. So $\pi(\langle\alpha,\beta\rangle)=\langle\pi(\alpha),\pi(\beta)\rangle\in g(a)$. By the definition, $\pi(\alpha)\Vdash\pi(\check{\beta})\in\dot{x}_{a}$. Hence $\pi(\beta)\in\dot{x}_{a}^{\pi[G_{2}*H_{2}\cap M]}$ and thus $\beta\in\dot{f}^{G_{2}*H_{2}}(a)$. Let $\beta\in\dot{f}^{G_{2}*H_{2}}(a)$. Then $\pi(\beta)\in\dot{x}_{a}^{\pi[G_{2}*H_{2}\cap M]}$. Ergo there exists $\pi(\alpha)\in\pi[G_{2}*H_{2}\cap M]$ such that $\pi(\alpha)\Vdash\pi(\beta)\in\dot{x}_{a}$. Hence $\pi(\langle\alpha,\beta\rangle)\in g(a)$ and $\langle\alpha,\beta\rangle\in e(a)=b_{e}\cap M$. Thus $\beta\in b_{f}\cap M$. ∎
@@ 8. Open Questions
We have shown a slight variation of the Iterands of the Cummings-Foreman iteration (see [2]) can force $\operatorname{ISP}(\tau^{+},\kappa,\geq\kappa)$ to be indestructible under ${<}\,\kappa$-directed closed forcing (which is an important step in showing that the tree property holds at all $\aleph_{n+2}$), hopefully making progress toward an answer to the questions.

### v2

@@ 7. Indestructibility of $\operatorname{ISP}$
We can view $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)*\dot{\mathbb{L}}$ as an ordering on a product $\mathbb{P}\times\mathbb{Q}$ by letting $\mathbb{P}:=\operatorname{Add}(\tau,\kappa\smallsetminus\xi)$ and $\mathbb{Q}$ consist of tripels $(q,r,\sigma)$ with $q,r$ as in $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)$ and $\sigma$ such that $\Vdash\sigma\in\dot{\mathbb{L}}$. This ordering is easily seen to be iteration-like. Moreover, the base ordering is $\tau^{+}$-Knaster (this is clear) and the term ordering is $\mu-$closed: Let $(p,(q_{\alpha},r_{\alpha},\sigma_{\alpha}))_{\alpha<\delta}$ for $\delta<\mu$ be a descending sequence. Then the sequence $(p,(q_{\alpha},r_{\alpha}))_{\alpha<\delta}$ is a descending sequence in the term ordering on $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)$. Therefore we can find a lower bound $(p,(q,r))$ of $(p,(q_{\alpha},r_{\alpha}))_{\alpha<\delta}$ in $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)$. Then $(p,(q,r))$, being a lower bound, forces the sequence $(\sigma_{\alpha})_{\alpha<\delta}$ to be descending in the $\kappa-$directed closed forcing $\dot{\mathbb{L}}$. Hence by the maximal principle we can fix $\sigma$ which is forced to be a lower bound of $(\sigma_{\alpha})_{\alpha<\delta}$. In summary, $(p,(q,r,\sigma))$ is a lower bound of $(p,(q_{\alpha},r_{\alpha},\sigma_{\alpha}))_{\alpha<\delta}$.
Now apply Theorem 3.8. ∎
Now we can show that $\mathbb{M}_{3}^{l}(\tau,\mu,\kappa)$ forces $\operatorname{ISP}$ in such a way that it is indestructible under $\kappa-$directed closed forcing.
@@
Let $\tau<\mu<\kappa$ be regular cardinals such that $\tau^{<\tau}=\tau$ and $\kappa$ is supercompact. Let $l$ be a Laver function. Then $\mathbb{M}_{3}^{l}(\tau,\mu,\kappa)$ forces $\operatorname{ISP}(\tau^{+},\kappa,\geq\kappa)$ and it is indestructible under $\kappa-$directed closed forcing.
@@
For any ordinal $\xi$, let $\mathbb{M}(\xi):=\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$.
It suffices to show that for any $\mathbb{M}(\kappa)$-name $\dot{\mathbb{L}}$ for a $\kappa-$directed closed forcing, $\mathbb{M}(\kappa)*\dot{\mathbb{L}}$ forces $\operatorname{ISP}(\tau^{+},\kappa,\geq\kappa)$. Additionally, it suffices to show that $\operatorname{ISP}(\tau^{+},\kappa,\lambda)$ is forced for arbitrarily large $\lambda$.
Thus let $\dot{\mathbb{L}}$ be an $\mathbb{M}(\kappa)$-name for a $\kappa-$directed closed forcing, let $\lambda\geq|\dot{\mathbb{L}}|$ and $\dot{f}$ an $\mathbb{M}(\kappa)*\dot{\mathbb{L}}$-name for a $\tau^{+}-$slender $(\kappa,\lambda)$-list, forced by some $(p,\sigma)\in\mathbb{M}(\kappa)*\dot{\mathbb{L}}$. Let $\dot{F}$ be an $\mathbb{M}(\kappa)*\dot{\mathbb{L}}$-name for a function such that any $M\prec H(\Theta^{\prime})$ closed under $\dot{F}$ witnesses the slenderness. We will view $\mathbb{M}(\kappa)*\dot{\mathbb{L}}$ as a partial order on $\lambda$. Let $\Theta$ be large and $M\prec H(\Theta)$ with the following:
@@
Then $p$ is actually a condition in $\mathbb{M}(\nu)$ and $\pi(\sigma)$ is an $\mathbb{M}(\nu)$-name for a condition in $\pi(\dot{\mathbb{L}})$. Thus we can view $(p,\pi(\sigma))$ as a condition $q$ in $\mathbb{M}(\nu+1)$. Let $G_{0}$ be an $\mathbb{M}(\kappa)$-generic filter containing $q$ and work in $V[G_{0}]$.
The collapse $\pi\colon M\to N$ extends to $M[G_{0}]\to N[\pi[G_{0}\cap M]]$ and, by the $\nu$-cc. of $\mathbb{M}(\nu)$, $G_{0}^{\prime}:=\pi[G_{0}\cap M]$ is $\mathbb{M}(\nu)$-generic over $V$. Let $H_{0}^{\prime}$ be the $\pi(\dot{\mathbb{L}})$-generic filter induced by $G$. Then $\pi^{-1}[H_{0}^{\prime}]\cup\{\sigma^{G}\}$ is (by elementarity) a $\kappa-$sized collection of pairwise compatible conditions in $\dot{\mathbb{L}}^{G}$. Ergo there exists a lower bound $r$. If now $H_{0}$ is $\dot{\mathbb{L}}^{G}$-generic containing $r$, $\pi[H_{0}\cap M]=H_{0}^{\prime}$ is $\pi(\dot{\mathbb{L}})$-generic over $V[G_{0}^{\prime}]$. This shows that for any model $M$ as above, there exists a condition $q_{M}\leq(p,\sigma)$ forcing that for any $\mathbb{M}(\kappa)*\dot{\mathbb{L}}$-generic filter GHG*H, $\pi[G*H\cap M]$ is equal to the $\pi(\mathbb{M}(\kappa)*\dot{\mathbb{L}})$-generic filter induced by $G$.
Now we transform, as before, $\dot{f}$ into a ground-model $(\kappa,\lambda)$-list. Let $a\in[\lambda]^{<\kappa}$.
@@
For simplicity we write $\pi_{M[G_{1}*H_{1}]}=\pi$.
Toward a contradiction, assume that $\pi[\dot{f}(\check{a})]\notin V[G_{1}^{\prime}*H_{1}^{\prime}]$. As $V[G_{1}*H_{1}]$ is an extension of $V[G_{1}^{\prime}*H_{1}^{\prime}]$ by a forcing with the $\tau^{+}-$approximation property, there exists $z\in[\pi(\lambda)]^{<\tau^{+}}\cap V[G_{1}^{\prime}*H_{1}^{\prime}]$ such that $z\cap\pi[\dot{f}^{G_{1}*H_{1}}(a)]\notin V[G_{1}^{\prime}*H_{1}^{\prime}]$. Because $H_{1}^{\prime}$ is generic for a $\nu-$directed closed forcing, $z\in V[G_{1}^{\prime}]$ and by Lemma 4.6, $z\in N[G_{1}^{\prime}]$, so $z=\pi(y)$ for $y\in M[G_{1}]\subseteq M[G_{1}*H_{1}]$. Furthermore, $M[G_{1}*H_{1}]\cap H(\Theta^{\prime})$ is closed under $\dot{F}^{G_{1}*H_{1}}$, so it witnesses the $\tau^{+}-$slenderness of $\dot{f}^{G_{1}*H_{1}}$. But then
@@
a contradiction. So $\pi[\dot{f}^{G_{1}*H_{1}}(a)]\in V[G_{1}^{\prime}*H_{1}^{\prime}]$ and there exists a condition $q_{a}$ as required.
Lastly, we want to show that $q_{a}$ forces that $M[\Gamma]$ is a $\lambda$-ineffability witness for $\kappa$ with respect to $\dot{f}$. To this end, let $G_{2}*H_{2}$ be $\mathbb{M}(\kappa)*\dot{\mathbb{L}}$-generic containing $q_{a}$. By assumption we are in case (1) and there is $b_{e}\in M$ such that $b_{e}\cap M=e(M\cap\lambda)$. Define
@@
where the last equality holds because $q_{a}\leq q_{M}$, which implies $M[G_{2}*H_{2}]\cap V=M$.
Let $\beta\in b_{f}\cap M$. By elementarity there is $\alpha\in G_{2}*H_{2}\cap M$ such that $\langle\alpha,\beta\rangle\in b_{e}$ and we have $\langle\alpha,\beta\rangle\in b_{e}\cap M=e(a)=\pi^{-1}(g(a))$. So $\pi(\langle\alpha,\beta\rangle)=\langle\pi(\alpha),\pi(\beta)\rangle\in g(a)$. By the definition, $\pi(\alpha)\Vdash\pi(\check{\beta})\in\dot{x}_{a}$. Hence $\pi(\beta)\in\dot{x}_{a}^{\pi[G_{2}*H_{2}\cap M]}$ and thus $\beta\in\dot{f}^{G_{2}*H_{2}}(a)$.
Let $\beta\in\dot{f}^{G_{2}*H_{2}}(a)$. Then $\pi(\beta)\in\dot{x}_{a}^{\pi[G_{2}*H_{2}\cap M]}$. Ergo there exists $\pi(\alpha)\in\pi[G_{2}*H_{2}\cap M]$ such that $\pi(\alpha)\Vdash\pi(\beta)\in\dot{x}_{a}$. Hence $\pi(\langle\alpha,\beta\rangle)\in g(a)$ and $\langle\alpha,\beta\rangle\in e(a)=b_{e}\cap M$. Thus $\beta\in b_{f}\cap M$. ∎
@@ 8. Open Questions
We have shown a slight variation of the iterands of the Cummings-Foreman iteration (see [3]) can force $\operatorname{ISP}(\tau^{+},\kappa,\geq\kappa)$ to be indestructible under $\kappa-$directed closed forcing (which is an important step in showing that the tree property holds at all $\aleph_{n+2}$), hopefully making progress toward an answer to the questions.